. A thermal neutron has a speed of ν that corresponds to room temperature T=300K. What is the wavelength of a thermal neutron ?


1
Expert's answer
2020-08-17T09:03:41-0400

Given

Temperature is "T=300 K"

Mass of neutron is "m=1.67\times10^{-27} kg"

The wavelength of thermal neutrons is

"\lambda=\frac{h}{\sqrt{3mk_BT} }"

"=\frac{6.626\times10^{-34}}{\sqrt{3(1.67\times10^{-27})(1.38\times10^{-23})(300)} }"

"=1.45\times10^{-10} m"